OpenStudy (anonymous):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

pooja195 (pooja195):



OpenStudy (anonymous):

45 and 2

pooja195 (pooja195):

2 is wrong... 5 x -3 remember we are multiplying not adding

OpenStudy (anonymous):

ok

pooja195 (pooja195):

So what is 5 x -3?

OpenStudy (anonymous):

srry afk -15

pooja195 (pooja195):

Good so we have \[\huge~\rm~ 45-15=w • 9 – 5 • 3 \] Now we need to multiply the other side \[\huge~\rm~9(w)-5(3)\] What do you get when you multiply those?

OpenStudy (anonymous):

9w - 15

pooja195 (pooja195):

Well done :) now we have \[\huge~\rm~45−15=9w-15\] Any ideas on what the next step may be?

OpenStudy (anonymous):

subtract 15 on both sides?

pooja195 (pooja195):

No we do the opposite of subtraction....remember we are trying to cancel them out

OpenStudy (anonymous):

30 - 15 = 9w

OpenStudy (anonymous):

oh ok

pooja195 (pooja195):

\[\huge~\rm~45−15+15=9w−15+15\]

OpenStudy (anonymous):

45 - 30 = 9w - 30

pooja195 (pooja195):

-15+15=? it is not -30

OpenStudy (anonymous):

oh ok 15 then

OpenStudy (anonymous):

oh i didnt say - 30 i just meant 30

pooja195 (pooja195):

Try again. if you have -15 and add 15 what do you get?

OpenStudy (anonymous):

yes sorry its 0

pooja195 (pooja195):

Yes so \[\huge~\rm~45=9w\] Well now that those cancel out what is the next step?

OpenStudy (anonymous):

its ok i learned it in 4th grade im just a little off track

OpenStudy (anonymous):

add 45 on both sides

pooja195 (pooja195):

Now we are trying to isolate the w right? So we need to divide both sides by 9w

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Let me know what you get as an answer

OpenStudy (anonymous):

-36 = w

pooja195 (pooja195):

The w part is correct but what is 45/9 ?

OpenStudy (anonymous):

5

pooja195 (pooja195):

Good so w=?

OpenStudy (anonymous):

i thought we were subtracting

OpenStudy (anonymous):

5?

pooja195 (pooja195):

Yes and we dont subtract here.

pooja195 (pooja195):

We divide

OpenStudy (anonymous):

ok so my answer is right

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

ok thx m8

pooja195 (pooja195):

:)